OpenStudy (anonymous):

Andrea is having her yard landscaped. She obtained an estimate from two landscaping companies. Company A gave an estimate of $220 for materials and equipment rental plus $65 per hour for labor. Company B gave an estimate of $300 for materials and equipment rental plus $55 per hour for labor. Determine how many hours of labor will be required for the two companies to cost the same. A) 8 hours B) 12 hours C) 11 hours D) 7 hours

OpenStudy (anonymous):

A(x)=200+65*x B(x)=300+55*x A(x)=B(x) 200+65*x=300+55*x 65*x-55*x=300-220 10x=80 x=8 A) 8 hours